In Math when you are dealing with Integers and Division, how do you divide with Like and Unlike signs?
Found 2 solutions by tutorcecilia, doukungfoo:
Answer by tutorcecilia(2152)

The rules are the same as multiplying with negative and positve numbers:
If the signs are the same, the answer will be positive.
Example: 6/3=2 and -6/-3=2
.
If the signs are different, the answer will be negative.
-6/3=-2 and 6/-3=-2

Answer by doukungfoo(195)

Dividing numbers with the same sign will result in a positive answer.
Example:
4/2 = 2
-4/-2 = 2
Dividing numbers with different signs will result in a negative answer.
Example:
-4/2 = -2
4/-2 = -2